### v3

@@ 4.2 Quantitative results
Next, we replace the above point prompts with text prompts corresponding to the objects. Given a text input, we prompt the Grounding DINO [18] to generate target bounding boxes, which serve as input prompts for the SAM to obtain 2D segmentation masks. Then, following the segmentation process in Section 3, we can transform the 2D masks into 3D Gaussian masks corresponding to the text input. Similar to the above evaluation process, we also calculate the IoU and Acc between 2D rendered masks and given ground-truth masks. Quantitative results are shown in Table 2. For both room and truck scenes, the IoU and Acc values can achieve a similar level as using clicked points as input prompts. These results demonstrate our method can combine multi-modal prompts as input.
4.3 Qualitative results
@@
Figure 5: 3D segmentation results with text prompts in Mip-NeRF360-garden [2].
@@
Text-Guided Segmentation: We further conduct 3D object segmentation experiments guided by text prompts. We follow the same process as in Table 2 experiments to conduct visualization experiments. In order to compare our method with SA3D under the same setting, we choose the same garden scene from Mip-NeRF 360 dataset [2], using three text prompts including “The table”, “The vase” and “The bonsai”. Results in Fig. 5 show that our approach can achieve accurate segmentation results by simply providing object names, demonstrating our method’s potential in combining with language models. Compared with the SA3D in the “The table” case, our approach can segment the complete object with desired table legs, while SA3D only gives the tabletop.
Comparison between 2D and 3D Segmentation: Our method improves the segmentation quality of SAM. Due to the sensitivity of SAM to the scene viewpoint, the 2D segmentation results might be incomplete under some views, but our 3D segmentation can still guarantee complete segmentation results via a multi-view ensemble (Fig. 6).
Table 3: Ablation on Gaussian decomposition (GD) on SPIn-NeRF dataset. The 2nd and 3rd columns compare the results before and after using the Gaussian decomposition (GD). The last column represents the results of directly removing the gaussians across mask boundaries.
@@
Pinecone 92.1 98.9 91.6 98.9 82.9 97.5
Truck 94.0 97.9 93.4 97.8 91.4 96.8
Lego 88.4 99.7 88.4 99.6 82.9 99.4
mean 89.9 98.7 86.6 98.1 84.2 97.7
Figure 6: Comparison between 2D and 3D Segmentation.
4.4 Ablation Study
Gaussian Decomposition:
Gaussian Decomposition is proposed to address the issue of roughness boundaries of 3D segmented objects, which results from the non-negligible spatial sizes of 3D Gaussian located at the boundary. We conduct both quantitative and qualitative experiments to verify the effectiveness of our approach. For quantitative ablation experiments, Table 3 shows the results. We compare our proposed Gaussian Decomposition scheme with two other processing methods, one for segmentation without any special handling (the 3rd column), and the other for directly removing such Gaussians (the last column). The experiments are conducted on SPIn-NeRF dataset, following the same evaluation process as Table 1. Compared with the others, our proposed approach outperforms them on all scenes with +3.3% and +5.7% IoU, respectively, demonstrating the effectiveness of Gaussian Decomposition. Besides, it can seen that directly removing these gaussians cannot decrease the roughness of mask boundaries, leading to even 2.4% lower IoU values.
Fig. 8 shows the visualization results for comparison. We select two representative scenes in which the junction/contact exists large-scale Gaussian points (can be seen in the 2nd column). This issue is alleviated after utilizing the proposed Gaussian Decomposition strategy. Though this idea is simple, the improvement is obvious.
Table 4: Ablation on different number of views for 3D segmentation. Numbers in parentheses represent the used view percentage of the total training view.
@@ 4.4 Ablation Study
IoU on truck 90.27 90.97 92.11 93.48
Table 5: Ablation on confidence score threshold $\tau$ on the “truck” scene.
$\tau$ 0.3 0.5 0.6 0.7 0.9
IoU 61.7 82.7 93.4 90.6 53.9
Selection of Views:
In this section, we study the influence of selected 2D views on 3D Gaussian segmentation. Fig. 9 (b) demonstrates the segmentation quality using different numbers of views. It is noteworthy that the leaves in the “fern” scene are very tiny and challenging. Even so, with only two sparse views, our approach can achieve decent results, and the segmentation quality quickly improves when increasing the number of views. Table 4 also draws a similar conclusion: as the number of views increases, the IoU values will also increase. It is worth noting that even with sparse views (10% percentage), we can obtain relatively decent results for both two types of scenes, which also demonstrates the robustness and effectiveness of our approach. In practical applications, using sparse views (below 10% of total views) will greatly improve efficiency, although under 100% views, our method can still be completed within one minute.
@@
Hyper-Parameters:
In the process of our method, we set a hyper-parameter $\tau$ to control the threshold of label confidence score in Sec. 3.5. As shown in Table 5, too large $\tau$ values will lead to the partially missing the segmented objects, and too small $\tau$ values will remain more background objects. For scenes with some views that may cover only part of the objects to be segmented (i.e. half of the truck body in the “truck” scene), a small $\tau$ value is preferred. In the interactive segmentation process, this threshold can be set manually according to the complexity of the scenes and the quality of camera views.
5 Discussions and Limitations
@@
Figure 9: Ablation on view selection.
6 Conclusion

### v4

@@ IV-D Boundary Segmentation Analysis
TABLE III: Evaluation on boundary masks. We extract parts of 3-pixel width alongside boundaries of ground truth and segmentation masks and evaluate on IoU, AP, and F1-score metrics.
Scene IoU AP F1-score
SAGA [6] Ours SAGA Ours SAGA Ours
pinecone 13.8 21.8 15.7 23.5 24.0 35.6
fortress 16.7 23.3 31.8 60.2 28.5 37.2
IV-E Qualitative results
@@
Figure 5: 3D segmentation with text prompts in Mip-NeRF360-garden [3].
@@
Text-Guided Segmentation: We further conduct 3D object segmentation experiments guided by text prompts. We follow the same process as in Table II experiments to conduct visualization experiments. In order to compare our method with SA3D under the same setting, we choose the same garden scene from Mip-NeRF 360 dataset [3], using three text prompts including “The table”, “The vase” and “The bonsai”. Results in Fig. 5 show that our approach can achieve accurate segmentation results by simply providing object names, demonstrating our method’s potential in combining with language models. Compared with the SA3D in the “The table” case, our approach can segment the complete object with desired table legs, while SA3D only gives the tabletop.
Comparison between 2D and 3D Segmentation: Our method improves the segmentation quality of SAM. Due to the sensitivity of SAM to the scene viewpoint (e.g. illumination), the 2D segmentation results might be incomplete under some views, but our 3D segmentation can still guarantee complete segmentation results via a multi-view ensemble (Fig. 7, $1^{st}$ Row). Furthermore, we can obtain detailed 3D segmentation results under noisy predictions from SAM. As shown in Fig. 7, $2^{nd}$ Row, the multi-view voting ensures that object Gaussians are retained correctly and background Gaussians obtained from the noisy boundaries of SAM predictions are removed.
IV-F Time Cost Analysis
We conduct the time cost analysis compared with previous 3DGS segmentation methods, namely SAGA [6] and Gaussian Grouping [92]. We select the figurines scene from the LERF dataset [36], which includes 20+ objects, using a single NVIDIA A100 GPU for fair evaluation. All methods utilize SAM to extract image features and then perform mask generation. Thus, they share a similar time cost, roughly 2 minutes to process 300 images (SAM Extraction in Tab. IV). This procedure can be regarded as a pre-processing and can be accelerated by image-batching or deploying Efficient-SAM. For extra processing time, both baseline methods require gradient descent optimization over 30,000 iterations to distill SAM features [6] or ID embedding [92] into 3D space linked with each 3D Gaussian, resulting in considerable additional training time for re-optimizing 3D scenes. On the contrary, our method only needs to perform multi-view and multi-label assignments, the time cost of which is merely 12 seconds. As for segmenting a single object, our method merely requires an efficient voting process, taking just 0.4 milliseconds, which is much faster than other methods as they both demand network forwarding.
Figure 7: Comparison between 2D and 3D Segmentation. Noisy predictions exist in SAM’s segmentation, such as incomplete objects and background regions. These issues are solved in our 3D segmentation.
TABLE IV: Time cost comparisions on figurines scene with 20+ objects to be segmented simultaneously.
Time SAM Extra Optimization Forward
Extraction Processing Steps Segmentation
SAGA [6] 2 min 10 min 30000 0.5s
Ga-Grouping [92] 9 min 30000 0.3s
Ours 12s 1 0.4ms
TABLE V: Ablation on Gaussian decomposition (GD) on SPIn-NeRF dataset. The 2nd and 3rd columns compare the results before and after using the Gaussian decomposition (GD). The last column represents the results of directly removing the gaussians across mask boundaries.
@@ IV-F Time Cost Analysis
Pinecone 92.6 98.9 91.6 98.9 82.9 97.5
Truck 93.0 97.9 93.4 97.8 91.4 96.8
Lego 90.2 99.7 88.4 99.6 82.9 99.4
mean 90.0 98.7 86.6 98.1 84.2 97.7
IV-G Ablation Study
Gaussian Decomposition: Gaussian Decomposition is proposed to address the issue of roughness boundaries of 3D segmented objects, which results from the non-negligible spatial sizes of 3D Gaussian located at the boundary. We conduct both quantitative and qualitative experiments to verify the effectiveness of our approach. For quantitative ablation experiments, Table V shows the results. We compare our proposed Gaussian Decomposition scheme with two other processing methods, one for segmentation without any special handling (the 3rd column), and the other for directly removing such Gaussians (the last column). The experiments are conducted on SPIn-NeRF dataset, following the same evaluation process as Table I. Compared with the others, our proposed approach outperforms them on all scenes with +3.4% and +5.8% IoU, respectively, demonstrating the effectiveness of Gaussian Decomposition. Besides, it can seen that directly removing these gaussians cannot decrease the roughness of mask boundaries, leading to even 2.4% lower IoU values.
Fig. 9 shows the visualization results for comparison. We select two representative scenes in which large-scale Gaussian points exist at the junction/contact (can be seen in the 2nd column). This issue is alleviated after utilizing the proposed Gaussian Decomposition strategy. Though this idea is simple, the improvement is obvious.
TABLE VI: Ablation on different number of views for 3D segmentation. Numbers in parentheses represent the used view percentage of the total training view.
@@ IV-G Ablation Study
IoU on truck 90.27 90.97 92.11 93.0
TABLE VII: Ablation on confidence score threshold $\tau$ on the truck scene.
Scene / $\tau$ 0.6 0.65 0.7 0.8
pinecone 90.8 92.0 92.6 91.4
fortress 94.0 95.4 95.6 96.2
lego 89.7 89.8 90.2 90.2
Figure 8: Ablation on view selection, including view directions and the number of views.
Selection of Views: In this section, we study the influence of selected 2D views on 3D Gaussian segmentation. Fig. 8 (b) demonstrates the segmentation quality using different numbers of views. It is noteworthy that the leaves in the “fern” scene are very tiny and challenging. Even so, with only two sparse views, our approach can achieve decent results, and the segmentation quality quickly improves when increasing the number of views. Table VI also draws a similar conclusion: as the number of views increases, the IoU values will also increase. It is worth noting that even with sparse views (10% percentage), we can obtain relatively decent results for both two types of scenes, which also demonstrates the robustness and effectiveness of our approach. In practical applications, using sparse views (below 10% of total views) will greatly improve efficiency, although under 100% views, our method can still be completed within one minute.
@@
Hyper-Parameters: In the process of our method, we set a hyper-parameter $\tau$ to control the threshold of label confidence score in Sec. III-E. To evaluate the robustness of the selection of the hyper-parameter under different scenes, we set different values to conduct experiments on Scene pinecone, fortress, and lego from LLFF [57] dataset. As indicated in Tab. VII, our method is not sensitive to the threshold selection. Robust results are observed in multiple scenes when setting the threshold between 0.6 and 0.8. Values within this range can obtain fine results. In all of our experiments, the value of the hyper-parameter $\tau$ is set to 0.7 by default.
V Discussions and Limitations
@@
VI Conclusion